a box with a square base and no top is to be made from a square piece of cardboard by cutting 4-in. squares from each corner and folding up the sides. the box is to hold 100 in.(cubed). how big a piece of cardboard is needed?
Answer by mananth(16949)

let the side of the square cardboard be x
4 in corners are to be cut
the dimensions of base will be (x-8)
Volume of Box = L*W*H
100=(x-8)(x-8)*4
/4
25=(x-8)^2
25=x^2-16x+64
x^2-16x+39=0
x^2-13x-3x+39=0
x(x-13)-3(x-13)=0
(x-13(x-3)=0
x=13 OR 3
3 is not possible
so 13 in is the length of side of carboard
m.ananth@hotmail.ca